OpenStudy (anonymous):

a right triangle has an angle that measures 34 and the adjacent side measure 17. what is the lenght of the hypotenuse to the nearset tenth?

OpenStudy (anonymous):

 \[\cos \theta = adjacent / hypotenuse\] so, \[hypotenuse = adjacent / \cos \theta\] Can you find the answer now?